# IndicMMLU-Pro: Benchmarking the Indic Large Language Models

### v1

@@ Abstract
Known by more than 1.5 billion people in the Indian subcontinent, Indic languages present unique challenges and opportunities for natural language processing (NLP) research due to their rich cultural heritage, linguistic diversity, and complex structures. IndicMMLU-Pro is a comprehensive benchmark designed to evaluate Large Language Models (LLMs) across Indic languages, building upon the MMLU Pro (Massive Multitask Language Understanding) framework. Covering major languages such as Hindi, Bengali, Gujarati, Marathi, Kannada, Punjabi, Tamil, Telugu, and Urdu, our benchmark addresses the unique challenges and opportunities presented by the linguistic diversity of the Indian subcontinent. This benchmark encompasses a wide range of tasks in language comprehension, reasoning, and generation, meticulously crafted to capture the intricacies of Indian languages. IndicMMLU-Pro provides a standardized evaluation framework to push the research boundaries in Indic language AI, facilitating the development of more accurate, efficient, and culturally sensitive models. This paper outlines the benchmarks’ design principles, task taxonomy, data collection methodology, and presents baseline results from state-of-the-art multilingual models. As a publicly available resource, IndicMMLU-Pro [^1] https://huggingface.co/datasets/LinguaLift/IndicMMLU-Pro is set to contribute significantly to advancements in Indic language-based technologies and serve as a valuable tool for the NLP community.
IndicMMLU-Pro: Benchmarking the Indic Large Language Models
@@
Figure 1: IndicMMLU-Pro Dataset Construction and Evaluation Pipeline. The diagram illustrates the end-to-end process of creating and validating the IndicMMLU-Pro dataset across nine the Indic languages. Starting with the English MMLU-Pro dataset, content is translated using IndicTrans2 (1B parameters) and undergoes rigorous quality assurance through back-translation and multiple metric evaluations (chrF++, BLEU, METEOR, TER, and SacreBLEU). Only translations meeting quality thresholds proceed to the final dataset. The workflow also shows the comprehensive evaluation process including expert proofreading involving 13 reviewers who assess semantic accuracy, fluency, and linguistic style. This systematic approach ensures the creation of a high-quality, multilingual benchmark dataset that maintains the integrity of the original MMLU-Pro while adapting to the linguistic nuances of the Indic languages.
@@ 1 Introduction
With over 1.5 billion speakers, the Indic languages constitute a substantial component of the world’s linguistic tapestry, showcasing the incredible diversity of the Indian subcontinent, where languages from the Indo-Aryan and Dravidian families have evolved over centuries, shaped by a shared cultural and historical context. As an integral part of daily life, these languages facilitate communication and accessibility in various aspects of society, including education, government, media, healthcare, and social services. Further, by developing language-specific benchmarks and conducting research on the Indic languages, we can increase the downstream impact on these languages on inclusion, ultimately leading to better language supportand enhanced accessibility for Deaf and hard-of-hearing communities across the region who use the Indicsign languages Ananthanarayana et al. (2021b). This, in turn, can enable greater social inclusion, improved education and employment opportunities, and more effective participation in civic life for sign language users.
@@
The disparity in Natural Language Processing research and resources across languages is striking with a survey of existing literature showing how resources and research on the Indic languages lagging KJ et al. (2024a). In their work, Joshi et al. (2020) revealed a stark imbalance, where a mere 28 percent of languages are considered "winners," while a staggering 88 percent are "left behind." This disparity is exemplified by the contrast between English and Bengali, languages with comparable speaker populations Lane (2019). Despite this, English dominates Bengali regarding available resources, with hundreds of times more visibility on platforms like the Linguistic Data Consortium, Wikipedia, and academic publication venues of significance.
@@
Indic languages have thus historically received less attention in the field of natural language processing (NLP) compared to more globally dominant languages Das et al. (2024). The disparity in NLP resources for the Indic languages has been attributed to their remarkable linguistic diversity, intricate morphology, and scarcity of annotated datasets Kakwani et al. (2020a); Marreddy et al. (2022). Nevertheless, driven by the increasing demand for NLP applications in the Indic languages and the prospect of groundbreaking technological innovations Ananthanarayana et al. (2021a), there is an urgent need for rigorous evaluation benchmarks to accurately assess the performance of AI models in this domain.
@@
To address this gap, we introduce IndicMMLU-Pro, a benchmark built upon the principles of the recently released MMLU-Pro Wang et al. (2024) by Tiger Labs. IndicMMLU-Pro adapts the robust multi-task principles to the unique context of the Indic languages, providing a comprehensive evaluation framework that assesses the linguistic understanding, reasoning abilities, and generative capabilities of AI models.
@@
This paper makes three key contributions. Primarily, we introduce IndicMMLU-Pro, a novel benchmark for evaluating AI models across a wide range of tasks and multiple the Indic languages. Next, the design principles, task taxonomy, and data collection methodology of IndicMMLU-Pro are presented in detail, ensuring that the benchmark accurately captures the complex linguistic and cultural characteristics of the Indic languages. Furthermore, we establish baseline results on IndicMMLU-Pro using state-of-the-art multilingual models, laying the groundwork for future research and development of the IndicAI models.
@@ 2 Methodology
To create IndicMMLU-Pro, a benchmark for the Indic languages equivalent to MMLU Pro, we adopted a process similar to prior work on the IndicMMLU, organizing our approach into two main steps: dataset creation and baseline benchmarking.
@@ 2.1 Dataset Creation
Our goal was to provide MMLU-Pro in nine Indic languages: Hindi, Bengali, Telugu, Marathi, Tamil, Gujarati, Urdu, Kannada and Punjabi. To achieve this, we leveraged IndicTrans2 Gala et al. (2023), a state-of-the-art machine translation model specifically designed for the Indic languages.
As shown in Figure1, we used the IndicTrans2 to convert the questions and corresponding options from the original English MMLU Pro dataset into each of the target languages. This approach allowed us to maintain the structure and content of the original benchmark while adapting it to the linguistic characteristics of the Indic languages.
@@ 2.3 Dataset Structure
The resulting IndicMMLU-Pro dataset maintains the same structure as the original MMLU-Pro dataset, with separate subsets for each of the nine Indic languages. The dataset retains the original categories and task types from MMLU Pro, thus ensuring identical usage and facilitating comprehensive evaluation across various domains and cognitive skills.
@@ 2.5 Baseline Benchmarking
To establish baseline performance metrics for the IndicMMLU-Pro benchmark, we evaluated state-of-the-art multilingual language models: GPT-4o, GPT-4o-mini, Llama-3.1-8B-Instruct, IndicBERT Kakwani et al. (2020a), IndicBART Dabre et al. (2021), RemBERT Chung et al. (2020), MuRIL Khanuja et al. (2021), and XLM-RoBERTa Conneau et al. (2020), Navarasa, Airavata Gala et al. (2024), OpenHathi, TamilLlama Balachandran (2023), and MahaMarathi. These models were chosen due to their demonstrated capabilities in handling multiple languages and their specific design for inclusion of the Indic languages in their training data.
The benchmarking process was conducted as follows:
  • Model Selection: To provide a benchmark of performance over the various languages contained in IndicMMLU-Pro, we leverage the following language models with the Indic language understanding capabilities.
    • IndicBERT and IndicBART: Specifically designed for the Indic languages, offering robust performance in this domain.
@@
    • MuRIL: A multilingual model with a focus on Indian languages, providing comprehensive coverage of the Indic languages.
@@
    • Navarasa, Airavata, OpenHathi, TamilLlama, and MahaMarathi: Models specifically designed or fine-tuned for the Indic languages, enhancing their performance in this domain.
@@
The results of this baseline benchmarking are presented in Table 1 of the Results section, showing the accuracy scores for each model across all nine languages. This benchmarking process provides a solid foundation for understanding the current capabilities of multilingual models in the Indic languages and sets a baseline for future research and improvements in this area.
@@ 2.6 Proofreading and Setup
We conducted a proofreading exercise on 9,000 sentence pairs (English and Indic), where 1,000 of these pairs were from the test split of each of the nine languages in the IndicMMLU-Pro dataset. These sentences pairs were stratified equally across the 14 categories and their respective 4 sources in each of the nine languages. These sentence pairs were shuffled randomly across all the Indic languages to eliminate any bias and ensure there was no preservation for each source and category after the selection process.
The three criteria that were defined for proofreading scores were:
  • Semantic Accuracy and Correctness: ensuring the translation conveys the exact meaning of the original text without errors, omissions or additions.
@@ 3 Results
Table 2: Comparison of language model performance across Indian languages, both Indo-Aryan (i.e., Hindi, Bengali, Punjabi, Marathi, Urdu, and Gujrati) and Dravidan (i.e., Telegu, Tamil, and Kannada). Scores shown for Llama 3.1, Navarasa, Airavata, OpenHathi, TamilLlama, and MahaMarathi models where available.
Table 1 and 2 present the accuracy scores (in percentages) of various pre-trained language models evaluated on the IndicMMLU-Pro benchmark. The benchmark covers nine major Indic languages: Bengali, Gujarati, Hindi, Kannada, Marathi, Punjabi, Tamil, Telugu, and Urdu. The scores reflect the models’ ability to handle diverse linguistic challenges specific to the Indic languages.
@@ 3.2.1 Cross-linguistic Analysis
  • Performance Gap: A significant performance gap exists between GPT-4o/GPT-4o mini and other models, highlighting the advantages of larger, more advanced models in handling the Indic language complexities.
@@
Table 3: Back-translation evaluation metrics for the IndicMMLU-Pro dataset for 3 the Indic languages.
@@ 3.3 Dataset Quality Assessment
In order to heuristically assert the quality of the Indic language datasets, we convert the dataset from the respective Indic language back to English again by leveraging IndicTrans2 as in Section 2.1. The English translations are then scored by the following metrics:
@@
Preliminary assessments of the other languages (Bengali, Punjabi, Kannada, Telugu, Urdu, and Marathi) suggest similar trends in translation quality, with chrF++ scores ranging from 73 to 79. Full metrics for these languages are being compiled and will be included in future publications.
@@ 3.4 Cosine Similarity Scores
We evaluated the semantic similarity between the Indic languages from the IndicMMLU-Pro dataset and the English MMLU-Pro dataset using the LaBSE (Language-agnostic BERT Sentence Embedding) model. It loads both datasets and processes all the samples from each. For each pair of corresponding questions, it generates embeddings for the questions and their multiple-choice options using the LaBSE model.
The script then calculates cosine similarity between these embeddings to measure how semantically similar the Indic languages and English versions are. It computes two main metrics: average question similarity, and average choice similarity. The question and choice similarities indicate how close the meanings are between the two languages. This analysis helps assess how accurately the Indic languages in the IndicMMLU-Pro represent the original English MMLU-Pro in terms of meaning and structure, which is crucial for ensuring the quality and consistency of multilingual datasets in machine learning and natural language processing tasks.
@@ 4 Related Work
The study of multilingual language models has been significantly advanced by efforts to create and utilize datasets that encompass a wide range of languages. the Indic languages, with their diverse scripts and rich morphological structures, present unique challenges and opportunities for multilingual NLP research.
Multilingual datasets have been crucial in developing language models capable of understanding and generating text in multiple languages. The WikiAnn dataset, for instance, provides a valuable resource for named entity recognition across many languages, including several indic languages Pan et al. (2017). The IndicNLP Corpus Kunchukuttan (2020), developed as part of the IndicNLP Library, provides extensive monolingual corpora for various indic languages. This dataset has been instrumental in training and evaluating models specifically designed for these languages. IndicCorp Joshi et al. (2022), a large-scale dataset of the Indic languages, and the IndicNLPSuiteKakwani et al. (2020b), which includes pre-trained models and linguistic resources, have further enhanced the development of multilingual NLP models. These resources have provided a solid foundation for various NLP tasks such as machine translation, text classification, and more.
IndicGLUEKakwani et al. (2020c) is a comprehensive benchmark designed to evaluate the performance of NLP models on a variety of tasks across multiple indic languages. It includes tasks such as sentiment analysis, natural language inference, and question answering, making it a valuable resource for assessing the capabilities of multilingual models
@@
Aggarwal et al. (2022) introduce IndicXNLI, a benchmark specifically designed to evaluate the natural language inference capabilities of multilingual models for Indian languages. This benchmark includes a diverse set of inference tasks, reflecting the linguistic richness and complexity of the Indic languages. IndicXNLI is crucial for advancing the understanding and processing of Indian languages in NLP, providing a standard for evaluating inference models in this context.
Kumar et al. (2022) develop the IndicNLG benchmark, which provides multilingual datasets for a variety of natural language generation (NLG) tasks in the Indic languages. This benchmark addresses the unique challenges posed by the linguistic diversity of the Indic languages, offering datasets for tasks such as machine translation, summarization, and text generation. IndicNLG is instrumental in improving the performance and capabilities of NLG models for the Indic languages, fostering advancements in multilingual NLP.
MMLU (Massively Multilingual Language Understanding) Liang et al. (2020); Hu et al. (2020) is a comprehensive benchmark designed to evaluate the performance of language models across multiple languages and a wide range of tasks. This benchmark has been instrumental in assessing the capabilities of multilingual models and has provided insights into their strengths and limitations

### v2

# IndicMMLU-Pro: Benchmarking Indic Large Language Models on Multi-Task Language Understanding

@@ Abstract
Known by more than 1.5 billion people in the Indian subcontinent, Indic languages present unique challenges and opportunities for natural language processing (NLP) research due to their rich cultural heritage, linguistic diversity, and complex structures. IndicMMLU-Pro is a comprehensive benchmark designed to evaluate Large Language Models (LLMs) across Indic languages, building upon the MMLU Pro (Massive Multitask Language Understanding) framework. Covering major languages such as Hindi, Bengali, Gujarati, Marathi, Kannada, Punjabi, Tamil, Telugu, and Urdu, our benchmark addresses the unique challenges and opportunities presented by the linguistic diversity of the Indian subcontinent. This benchmark encompasses a wide range of tasks in language comprehension, reasoning, and generation, meticulously crafted to capture the intricacies of Indian languages. IndicMMLU-Pro provides a standardized evaluation framework to push the research boundaries in Indic language AI, facilitating the development of more accurate, efficient, and culturally sensitive models. This paper outlines the benchmarks’ design principles, task taxonomy, and data collection methodology, and presents baseline results from state-of-the-art multilingual models. As a publicly available resource, IndicMMLU-Pro [^1] https://huggingface.co/datasets/LinguaLift/IndicMMLU-Pro is set to contribute significantly to advancements in Indic language-based technologies and serve as a valuable tool for the NLP community.
IndicMMLU-Pro: Benchmarking Indic Large Language Models on Multi-Task Language Understanding
@@
Figure 1: IndicMMLU-Pro Dataset Construction and Evaluation Pipeline. The diagram illustrates the end-to-end process of creating and validating the IndicMMLU-Pro dataset across nine Indic languages. Starting with the English MMLU-Pro dataset, content is translated using IndicTrans2 (1B parameters) and undergoes rigorous quality assurance through back-translation and multiple metric evaluations (chrF++, BLEU, METEOR, TER, and SacreBLEU). Only translations meeting quality thresholds proceed to the final dataset. The workflow also shows the comprehensive evaluation process including expert proofreading involving 13 reviewers who assess semantic accuracy, fluency, and linguistic style. This systematic approach ensures the creation of a high-quality, multilingual benchmark dataset that maintains the integrity of the original MMLU-Pro while adapting to the linguistic nuances of Indic languages.
@@ 1 Introduction
With over 1.5 billion speakers, Indic languages constitute a substantial component of the world’s linguistic tapestry, showcasing the incredible diversity of the Indian subcontinent, where languages from the Indo-Aryan and Dravidian families have evolved over centuries, shaped by a shared cultural and historical context. As an integral part of daily life, these languages facilitate communication and accessibility in various aspects of society, including education, government, media, healthcare, and social services. Further, by developing language-specific benchmarks and conducting research on Indic languages, we can increase the downstream impact on these languages on inclusion, ultimately leading to better language support and enhanced accessibility for the deaf and hard-of-hearing communities across the region who use Indic sign languages Ananthanarayana et al. (2021b). This, in turn, can enable greater social inclusion, improved education and employment opportunities, and more effective participation in civic life for sign language users.
@@
The disparity in Natural Language Processing (NLP) research and resources across languages is striking with a survey of existing literature showing how resources and research on Indic languages lagging KJ et al. (2024a). In their work, Joshi et al. (2020) revealed a stark imbalance, where a mere 28 percent of languages are considered "winners," while a staggering 88 percent are "left behind." This disparity is exemplified by the contrast between English and Bengali, languages with comparable speaker populations Lane (2019). Despite this, English dominates Bengali regarding available resources, with hundreds of times more visibility on platforms like the Linguistic Data Consortium, Wikipedia, and academic publication venues of significance.
@@
Thus, Indic languages have historically received less attention in the field of NLP compared to more globally dominant languages Das et al. (2024). The disparity in NLP resources for Indic languages has been attributed to their remarkable linguistic diversity, intricate morphology, and the scarcity of annotated datasets Kakwani et al. (2020a); Marreddy et al. (2022). Nevertheless, driven by the increasing demand for NLP applications in Indic languages and the prospect of groundbreaking technological innovations Ananthanarayana et al. (2021a), there is an urgent need for rigorous evaluation benchmarks to assess the performance of AI models in this domain accurately.
@@
To address this gap, we introduce IndicMMLU-Pro, a benchmark built upon the principles of the recently released MMLU-Pro Wang et al. (2024) by Tiger Labs. IndicMMLU-Pro adapts the robust multi-task principles to the unique context of Indic languages, providing a comprehensive evaluation framework that assesses the linguistic understanding, reasoning abilities, and generative capabilities of AI models.
@@
This paper makes three key contributions. Primarily, we introduce IndicMMLU-Pro, a novel benchmark for evaluating AI models across a wide range of tasks and multiple Indic languages. Next, the design principles, task taxonomy, and data collection methodology of IndicMMLU-Pro are presented in detail, to ensure that the benchmark accurately captures the complex linguistic and cultural characteristics of Indic languages. Furthermore, we establish baseline results on IndicMMLU-Pro using state-of-the-art multilingual models, laying the groundwork for future research and development of Indic languages-based AI models.
@@ 2 Methodology
To create IndicMMLU-Pro, a benchmark for Indic languages equivalent to MMLU Pro, we adopted a process similar to the prior work on IndicMMLU, organizing our approach into two main steps: dataset creation and baseline benchmarking.
@@ 2.1 Dataset Creation
Our goal was to provide MMLU-Pro in nine Indic languages: Hindi, Bengali, Telugu, Marathi, Tamil, Gujarati, Urdu, Kannada and Punjabi. To achieve this, we used IndicTrans2 Gala et al. (2023), a state-of-the-art machine translation model specifically designed for Indic languages.
As shown in Figure1, we use IndicTrans2 to convert the questions and corresponding options from the original English MMLU Pro dataset into each of the target languages. This approach allowed us to maintain the structure and content of the original benchmark while adapting it to the linguistic characteristics of Indic languages.
@@ 2.3 Dataset Structure
The resulting IndicMMLU-Pro dataset maintains the same structure as the original MMLU-Pro dataset, with separate subsets for each of the nine Indic languages. The dataset retains the original categories and task types from MMLU Pro, thus ensuring identical usage and facilitating comprehensive evaluation in various domains and cognitive skills.
@@ 2.5 Baseline Benchmarking
To establish baseline performance metrics for the IndicMMLU-Pro benchmark, we evaluated state-of-the-art multilingual language models: GPT-4o, GPT-4o-mini, Llama-3.1-8B-Instruct, IndicBERT Kakwani et al. (2020a), IndicBART Dabre et al. (2021), RemBERT Chung et al. (2020), MuRIL Khanuja et al. (2021), and XLM-RoBERTa Conneau et al. (2020), Navarasa, Airavata Gala et al. (2024), OpenHathi, TamilLlama Balachandran (2023), and MahaMarathi. These models were chosen due to their demonstrated capabilities in handling multiple languages and their specific design for the inclusion of Indic languages in their training data. The benchmarking process was conducted as follows:
  • Model Selection: To provide a benchmark of performance over the various languages contained in IndicMMLU-Pro, we leverage the following language models with Indic language understanding capabilities.
    • IndicBERT and IndicBART: Specifically designed for Indic languages, offering robust performance in this domain.
@@
    • MuRIL: A multilingual model with a focus on Indian languages, providing comprehensive coverage of Indic languages.
@@
    • Navarasa, Airavata, OpenHathi, TamilLlama, and MahaMarathi: Models specifically designed or fine-tuned for Indic languages, enhancing their performance in this domain.
@@
The results of this baseline benchmarking are presented in Table 1 of the Results section, showing the accuracy scores for each model across all nine languages. This benchmarking process provides a solid foundation for understanding the current capabilities of multilingual models in Indic languages and sets a baseline for future research and improvements in this area.
@@ 2.6 Proofreading and Setup
We conducted a proofreading exercise on 9,000 sentence pairs (English and Indic), where 1,000 of these pairs were from the test split of each of the nine languages in the IndicMMLU-Pro dataset. These sentences pairs were stratified equally across the 14 categories and their respective 4 sources in each of the nine languages. These sentence pairs were shuffled randomly across all Indic languages to eliminate any bias and ensure there was no preservation for each source and category after the selection process. The three criteria that were defined for proofreading scores were:
  • Semantic Accuracy and Correctness: ensuring the translation conveys the exact meaning of the original text without errors, omissions, or additions.
@@ 3 Results
Table 2: Comparison of language model performance across Indian languages, both Indo-Aryan (i.e., Hindi, Bengali, Punjabi, Marathi, Urdu, and Gujarati) and Dravidan (i.e., Telegu, Tamil, and Kannada). Scores are shown for Llama 3.1, Navarasa, Airavata, OpenHathi, TamilLlama, and MahaMarathi models where available.
Table 1 and 2 present the accuracy scores (in percentages) of various pre-trained language models evaluated on the IndicMMLU-Pro benchmark. The benchmark covers nine major Indic languages: Bengali, Gujarati, Hindi, Kannada, Marathi, Punjabi, Tamil, Telugu, and Urdu. The scores reflect the models’ ability to handle diverse linguistic challenges specific to Indic languages.
@@ 3.2.1 Cross-linguistic Analysis
  • Performance Gap: A significant performance gap exists between GPT-4o/GPT-4o mini and other models, highlighting the advantages of larger, more advanced models in handling the complexities of Indic languages.
@@
Table 3: Back-translation evaluation metrics for the IndicMMLU-Pro dataset for 3 Indic languages.
@@ 3.3 Dataset Quality Assessment
In order to heuristically assert the quality of Indic language datasets, we convert the dataset from the respective Indic language back to English again by leveraging IndicTrans2 as in Section 2.1. The English translations are then scored by the following metrics:
@@
Preliminary assessments of the other languages (Bengali, Punjabi, Kannada, Telugu, Urdu, and Marathi) suggest similar trends in translation quality, with chrF++ scores ranging from 73 to 79. Complete metrics for these languages are being compiled and will be included in future publications.
@@ 3.4 Cosine Similarity Scores
We evaluated the semantic similarity between Indic languages from the IndicMMLU-Pro dataset and the English MMLU-Pro dataset using the LaBSE (Language-agnostic BERT Sentence Embedding) model. It loads both datasets and processes all the samples from each. For each pair of corresponding questions, it generates embeddings for the questions and their multiple-choice options using the LaBSE model.
The script then calculates cosine similarity between these embeddings to measure how semantically similar Indic languages and their english versions are. It computes two main metrics: average question similarity, and average choice similarity. The question and choice similarities indicate how close the meanings are between the two languages. This analysis helps assess how accurately Indic languages in the IndicMMLU-Pro represent the original English MMLU-Pro in terms of meaning and structure, which is crucial for ensuring the quality and consistency of multilingual datasets in machine learning and natural language processing tasks.
@@ 4 Related Work
The study of multilingual language models has been significantly advanced by efforts to create and utilize datasets that encompass a wide range of languages. Indic languages, with their diverse scripts and rich morphological structures, present unique challenges and opportunities for multilingual NLP research.
Multilingual datasets have been crucial in developing language models capable of understanding and generating text in multiple languages. The WikiAnn dataset, for instance, provides a valuable resource for named entity recognition across many languages, including several Indic languages Pan et al. (2017). The IndicNLP Corpus Kunchukuttan (2020), developed as part of the IndicNLP Library, provides extensive monolingual corpora for various Indic languages. This dataset has been instrumental in training and evaluating models specifically designed for these languages. IndicCorp Joshi et al. (2022), a large-scale dataset of Indic languages, and the IndicNLPSuiteKakwani et al. (2020b), which includes pre-trained models and linguistic resources, have further enhanced the development of multilingual NLP models. These resources have provided a solid foundation for various NLP tasks such as machine translation, text classification, etc.
IndicGLUEKakwani et al. (2020c) is a comprehensive benchmark designed to evaluate the performance of NLP models on a variety of tasks across multiple Indic languages. It includes tasks such as sentiment analysis, natural language inference, and question answering, making it a valuable resource for assessing the capabilities of multilingual models.
@@
Aggarwal et al. (2022) introduce IndicXNLI, a benchmark specifically designed to evaluate the natural language inference capabilities of multilingual models for Indian languages. This benchmark includes a diverse set of inference tasks, reflecting the linguistic richness and complexity of Indic languages. IndicXNLI is crucial for advancing the understanding and processing of Indian languages in NLP, providing a standard for evaluating inference models in this context.
Kumar et al. (2022) develop the IndicNLG benchmark, which provides multilingual datasets for a variety of natural language generation (NLG) tasks in Indic languages. This benchmark addresses the unique challenges posed by the linguistic diversity of Indic languages, offering datasets for tasks such as machine translation, summarization, and text generation. IndicNLG is instrumental in improving the performance and capabilities of NLG models for Indic languages, fostering advancements in multilingual NLP.
MMLU (Massively Multilingual Language Understanding) Liang et al. (2020); Hu et al. (2020) is a comprehensive benchmark designed to evaluate the performance of language models across multiple languages and a wide range of tasks. This benchmark has been instrumental in assessing the capabilities of multilingual models and has provided insights into their strengths and limitations.